OpenStudy (anonymous):

How much including VAT , did the stereo cost before the sale? The price of stereo is 120, & vat is 17 1/2, & the sale is 15 % off.

OpenStudy (anonymous):

\[120\times 1.175\times .85\]

OpenStudy (anonymous):

it says before sale, 120*1.175=141

OpenStudy (rulnick):

$120 / (100% - 15%) = $120 / 0.85 = $141.18 gives the original price before the sale, rounded to the nearest cent. (Check it: $141.18 at 15% off is $120.00.) So $141.18 (100% + 17.5%) = $141.18 (1.175) = $165.89 gives the price with VAT before the sale Answer: $165.89.